# Reality conditions for the KdV equation and quasi-periodic solutions in finite phase spaces

### 18 Dec 2023 v1

@@ Abstract.
In this paper reality conditions for finite-gap solutions of the KdV equation are determined. This new result brings the problem of quasi-periodic solutions of the KdV equation to a close. An effective computation of such solutions is suggested, and illustrated in genera $2$ and $3$.
Real-valued quasi-periodic solutions, which describe nonlinear waves, are proposed in every finite phase space of the hierarchy of hamiltonian systems of the KdV equation. These solutions are expressed in terms of the abelian $\wp_{1,1}$-function, which comes as a result of algebro-geometric integration. The construction of the hierarchy on coadjoint orbits of an $\mathrm{SL}(2)$ loop group is also addressed.
@@ 7.3. Genus 2
where $\omega^{\prime}_{k}=(\omega^{\prime}_{1,k},\omega^{\prime}_{3,k})^{t}$. First indices of the entries $\omega^{\prime}_{2i-1,k}$ correspond to the labels of holomorphic differentials $\mathrm{d}u_{2i-1}$, $i=1$, $2$, and $\omega^{\prime}_{1}$, $\omega^{\prime}_{2}$ are two columns of the period matrix $\omega^{\prime}$. On the subspace $-(\mathrm{b}x,\mathrm{b}t)^{t}+\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}$ the function $\wp_{1,1}$ is real-valued and bounded, with the critical values:
@@
Let $\mathcal{M}_{2}^{\circ}$ be fixed by $14\mathrm{b}^{2}=$ 1, $c_{4}=0$, $c_{3}=-15$, $c_{2}=-20$. Let the hamiltonians $h_{1}$, $h_{0}$ attains five values with different mutual positions of branch points: $h_{1}=-3$, $h_{0}=1$ $e_{1}\approx-6.99$, $e_{2}\approx-1.13$, $e_{3}\approx-0.39$ $e_{4}\approx 0.15$, $e_{5}\approx 8.35$; [Uncaptioned image] $h_{1}=100$, $h_{0}=-56$ $e_{1}\approx-5.69$, $e_{2}\approx-4.55$, $e_{3}\approx 0.72$ $e_{4}\approx 1.50$, $e_{5}\approx 8.01$; [Uncaptioned image] $h_{1}=100$, $h_{0}=81$ $e_{1}\approx-6.15$, $e_{2}\approx-3.52$, $e_{3}\approx-0.76$ $e_{4}\approx 2.48$, $e_{5}\approx 7.94$; [Uncaptioned image] $h_{1}=100$, $h_{0}=156$ $e_{1}\approx-6.30$, $e_{2}\approx-2.71$, $e_{3}\approx-1.67$ $e_{4}\approx 2.78$, $e_{5}\approx 7.90$; [Uncaptioned image] $h_{1}=210$, $h_{0}=460$ $e_{1}\approx-5.05$, $e_{2}\approx-3.86$, $e_{3}\approx-2.83$ $e_{4}\approx 4.76$, $e_{5}\approx 6.99$. [Uncaptioned image]
@@ 7.4. Genus $3$
$\begin{split}&\qquad C_{1}=\tfrac{1}{2}\omega^{\prime}_{1,1}+\tfrac{1}{2}\omega^{\prime}_{1,2}+\tfrac{1}{2}\omega^{\prime}_{1,3},\\ &\qquad C_{3}=\tfrac{1}{2}\omega^{\prime}_{3,1}+\tfrac{1}{2}\omega^{\prime}_{3,2}+\tfrac{1}{2}\omega^{\prime}_{3,3},\\ &\qquad C_{5}=-\mathrm{c}+\tfrac{1}{2}\omega^{\prime}_{5,1}+\tfrac{1}{2}\omega^{\prime}_{5,2}+\tfrac{1}{2}\omega^{\prime}_{5,3},\quad\mathrm{c}\in\operatorname{\mathbb{R}}.\end{split}$
@@
where $\omega^{\prime}_{k}=(\omega^{\prime}_{1,k},\omega^{\prime}_{3,k},\omega^{\prime}_{5,k})^{t}$. Recall that $\omega^{\prime}_{1}$, $\omega^{\prime}_{2}$, $\omega^{\prime}_{3}$ are three columns of the period matrix $\omega^{\prime}$. On $\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}+\tfrac{1}{2}\omega^{\prime}_{3}+\operatorname{\mathrm{Re}}\bar{\operatorname{\mathfrak{J}}}$ the function $\wp_{1,1}$ is real-valued and bounded, with the critical values:
$\begin{split}&\wp_{1,1}(\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}+\tfrac{1}{2}\omega^{\prime}_{3})=e_{1}+e_{4}+e_{5},\\ &\wp_{1,1}(\tfrac{1}{2}\omega_{1}+\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}+\tfrac{1}{2}\omega^{\prime}_{3})=e_{2}+e_{4}+e_{5},\\ &\wp_{1,1}(\tfrac{1}{2}\omega_{2}+\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}+\tfrac{1}{2}\omega^{\prime}_{3})=e_{1}+e_{3}+e_{5},\\ &\wp_{1,1}(\tfrac{1}{2}\omega_{1}+\tfrac{1}{2}\omega_{2}+\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}+\tfrac{1}{2}\omega^{\prime}_{3})=e_{2}+e_{3}+e_{5},\\ &\wp_{1,1}(\tfrac{1}{2}\omega_{3}+\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}+\tfrac{1}{2}\omega^{\prime}_{3})=e_{1}+e_{4}+e_{6},\\ &\wp_{1,1}(\tfrac{1}{2}\omega_{1}+\tfrac{1}{2}\omega_{3}+\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}+\tfrac{1}{2}\omega^{\prime}_{3})=e_{2}+e_{4}+e_{6},\\ &\wp_{1,1}(\tfrac{1}{2}\omega_{2}+\tfrac{1}{2}\omega_{3}+\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}+\tfrac{1}{2}\omega^{\prime}_{3})=e_{1}+e_{3}+e_{6},\\ &\wp_{1,1}(\tfrac{1}{2}\omega_{1}+\tfrac{1}{2}\omega_{2}+\tfrac{1}{2}\omega_{3}+\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}+\tfrac{1}{2}\omega^{\prime}_{3})=e_{2}+e_{3}+e_{6}.\end{split}$
@@
However, on the subspace $-(\mathrm{b}x,\mathrm{b}t,\mathrm{c})^{t}+\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}+\tfrac{1}{2}\omega^{\prime}_{3}$, which serves as the domain of the quasi-periodic KdV solution $\beta$, not more than one can be reachable, depending on the value of $\mathrm{c}$.
Plots to be added.

### 31 Dec 2023 v2

@@ Abstract.
In this paper reality conditions for finite-gap solutions of the KdV equation are determined. This new result brings the last piece of the puzzle of obtaining quasi-periodic solutions of completely integrable dynamical equations. Effective computation of the solutions is suggested, and illustrated in genera $2$ and $3$.
Real-valued quasi-periodic solutions, which describe nonlinear waves, are proposed in every finite phase space of the hierarchy of hamiltonian systems of the KdV equation. These solutions are expressed in terms of the abelian $\wp_{1,1}$-function, which comes as a result of algebro-geometric integration. The construction of the hierarchy on coadjoint orbits of an $\mathrm{SL}(2,\operatorname{\mathbb{R}})$ loop group is also addressed.
@@ 7.3. Genus 2
where $(\omega^{\prime}_{1,k},\omega^{\prime}_{3,k})^{t}=\omega^{\prime}_{k}$. First indices of the entries $\omega^{\prime}_{2i-1,k}$ correspond to the labels of holomorphic differentials $\mathrm{d}u_{2i-1}$, $i=1$, $2$, and $\omega^{\prime}_{1}$, $\omega^{\prime}_{2}$ are two columns of the period matrix $\omega^{\prime}$. On the subspace $-(\mathrm{b}x,\mathrm{b}t)^{t}+\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}$ the function $\wp_{1,1}$ is real-valued and bounded, with the critical values:
@@
Let $\mathcal{M}_{2}^{\circ}$ be fixed by $14\mathrm{b}^{2}=$ 1, $c_{4}=0$, $c_{3}=-15$, $c_{2}=-20$. Let the hamiltonians $h_{1}$, $h_{0}$ attains five values with different mutual positions of branch points: $h_{1}=-3$, $h_{0}=1$ $e_{1}\approx-6.99$, $e_{2}\approx-1.13$, $e_{3}\approx-0.39$ $e_{4}\approx 0.15$, $e_{5}\approx 8.35$; [Uncaptioned image]
$h_{1}=100$, $h_{0}=-56$ $e_{1}\approx-5.69$, $e_{2}\approx-4.55$, $e_{3}\approx 0.72$ $e_{4}\approx 1.50$, $e_{5}\approx 8.01$; [Uncaptioned image]
$h_{1}=100$, $h_{0}=81$ $e_{1}\approx-6.15$, $e_{2}\approx-3.52$, $e_{3}\approx-0.76$ $e_{4}\approx 2.48$, $e_{5}\approx 7.94$; [Uncaptioned image] $h_{1}=100$, $h_{0}=156$ $e_{1}\approx-6.30$, $e_{2}\approx-2.71$, $e_{3}\approx-1.67$ $e_{4}\approx 2.78$, $e_{5}\approx 7.90$; [Uncaptioned image] $h_{1}=210$, $h_{0}=460$ $e_{1}\approx-5.05$, $e_{2}\approx-3.86$, $e_{3}\approx-2.83$ $e_{4}\approx 4.76$, $e_{5}\approx 6.99$. [Uncaptioned image]
@@ 7.4. Genus $3$
$\begin{split}&\qquad C_{1}=\tfrac{1}{2}\omega^{\prime}_{1,1}+\tfrac{1}{2}\omega^{\prime}_{1,2}+\tfrac{1}{2}\omega^{\prime}_{1,3},\\ &\qquad C_{3}=\tfrac{1}{2}\omega^{\prime}_{3,1}+\tfrac{1}{2}\omega^{\prime}_{3,2}+\tfrac{1}{2}\omega^{\prime}_{3,3},\\ &\qquad C_{5}=-\mathrm{c}+\tfrac{1}{2}\omega^{\prime}_{5,1}+\tfrac{1}{2}\omega^{\prime}_{5,2}+\tfrac{1}{2}\omega^{\prime}_{5,3},\quad\mathrm{c}\in\operatorname{\mathbb{R}},\end{split}$
@@
where $(\omega^{\prime}_{1,k},\omega^{\prime}_{3,k},\omega^{\prime}_{5,k})^{t}=\omega^{\prime}_{k}$. Recall that $\omega^{\prime}_{1}$, $\omega^{\prime}_{2}$, $\omega^{\prime}_{3}$ are three columns of the period matrix $\omega^{\prime}$. On $\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}+\tfrac{1}{2}\omega^{\prime}_{3}+\operatorname{\mathrm{Re}}\bar{\operatorname{\mathfrak{J}}}$ the function $\wp_{1,1}$ is real-valued and bounded, with the critical values:
(88) $\begin{split}&\wp_{1,1}(\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}+\tfrac{1}{2}\omega^{\prime}_{3})=e_{1}+e_{4}+e_{5},\\ &\wp_{1,1}(\tfrac{1}{2}\omega_{1}+\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}+\tfrac{1}{2}\omega^{\prime}_{3})=e_{2}+e_{4}+e_{5},\\ &\wp_{1,1}(\tfrac{1}{2}\omega_{2}+\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}+\tfrac{1}{2}\omega^{\prime}_{3})=e_{1}+e_{3}+e_{5},\\ &\wp_{1,1}(\tfrac{1}{2}\omega_{1}+\tfrac{1}{2}\omega_{2}+\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}+\tfrac{1}{2}\omega^{\prime}_{3})=e_{2}+e_{3}+e_{5},\\ &\wp_{1,1}(\tfrac{1}{2}\omega_{3}+\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}+\tfrac{1}{2}\omega^{\prime}_{3})=e_{1}+e_{4}+e_{6},\\ &\wp_{1,1}(\tfrac{1}{2}\omega_{1}+\tfrac{1}{2}\omega_{3}+\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}+\tfrac{1}{2}\omega^{\prime}_{3})=e_{2}+e_{4}+e_{6},\\ &\wp_{1,1}(\tfrac{1}{2}\omega_{2}+\tfrac{1}{2}\omega_{3}+\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}+\tfrac{1}{2}\omega^{\prime}_{3})=e_{1}+e_{3}+e_{6},\\ &\wp_{1,1}(\tfrac{1}{2}\omega_{1}+\tfrac{1}{2}\omega_{2}+\tfrac{1}{2}\omega_{3}+\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}+\tfrac{1}{2}\omega^{\prime}_{3})=e_{2}+e_{3}+e_{6}.\end{split}$
@@
However, projection to the subspace $-(\mathrm{b}x,\mathrm{b}t,\mathrm{c})^{t}+\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}+\tfrac{1}{2}\omega^{\prime}_{3}$, which serves as the domain of the quasi-periodic KdV solution $\beta$, could contain not more than one of the values (88), if $\mathrm{c}$ is one of half-periods constructed from $\tfrac{1}{2}\omega_{1}$, $\tfrac{1}{2}\omega_{2}$, $\tfrac{1}{2}\omega_{3}$.
Let $\mathcal{M}_{3}^{\circ}$ be fixed by $14\mathrm{b}^{2}=$ 1, $c_{6}=0$, $c_{5}=-84$, $c_{4}=-160$, $c_{3}=7250$. Let the hamiltonians $h_{2}$, $h_{1}$, $h_{0}$ attain six values with different mutual positions of branch points. Computed by (87) with $\mathrm{b}=1/2$ and $\mathrm{c}=0$, shapes of $\beta(x,0)$, $x\in[0,-20(\omega_{1,1}+\omega_{1,3})]$, and $\beta(0,t)$, $t\in[0,10]$ or $t\in[0,20]$, are presented below.
[…]
$h_{2}=16678$, $h_{1}=-148832$, $h_{0}=-338529;\qquad$ $e_{1}\approx-11.44$, $e_{2}\approx-11.42$, $e_{3}\approx-5.67$ $e_{4}\approx-2.26$, $e_{5}\approx 5.41$, $e_{6}\approx 9.33$, $e_{7}\approx 16.04$;     [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$h_{2}=12678$, $h_{1}=-172935$, $h_{0}=-12782;\qquad$ $e_{1}\approx-11.95$, $e_{2}\approx-9.07$, $e_{3}\approx-8.97$ $e_{4}\approx-0.07$, $e_{5}\approx 4.98$, $e_{6}\approx 8.84$, $e_{7}\approx 16.23$;     [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$h_{2}=21460$, $h_{1}=-120322$, $h_{0}=-287405;\qquad$ $e_{1}\approx-12.88$, $e_{2}\approx-9.31$, $e_{3}\approx-6.21$ $e_{4}\approx-2.16$, $e_{5}\approx 4.31$, $e_{6}\approx 10.58$, $e_{7}\approx 15.67$;     [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$h_{2}=22460$, $h_{1}=-81268$, $h_{0}=-267380\qquad$ $e_{1}\approx-12.21$, $e_{2}\approx-10.94$, $e_{3}\approx-3.57$ $e_{4}\approx-3.56$, $e_{5}\approx 3.62$, $e_{6}\approx 11.23$, $e_{7}\approx 15.44$;     [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$h_{2}=32678$, $h_{1}=-63757$, $h_{0}=-372539\qquad$ $e_{1}\approx-13.85$, $e_{2}\approx-6.64$, $e_{3}\approx-5.10$ $e_{4}\approx-5.08$, $e_{5}\approx 3.35$, $e_{6}\approx 13.60$, $e_{7}\approx 13.72$.     [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
On fig. 6 shapes of $\beta(x,t)$, $x\in[0,6]$, $t\in[0,0.6]$ at some chosen values of $h_{2}$, $h_{1}$, $h_{0}$ are presented.
Figure 6. $N=3$. Quasi-periodic waves $\beta(x,t)$, $x\in[0,6]$, $t\in[0,0.6]$.